# LT3SD: Latent Trees for 3D Scene Diffusion

### v1

@@ 2 Related Work
3D Diffusion Models for Shape Generation. Diffusion models have shown unprecedented capability in generating 2D images and videos [44, 5]. Existing 3D diffusion work has focused on synthesizing 3D shapes [6, 18, 46, 56, 14, 30, 13, 20, 10, 52], showing quite promising results. Following LDMs [44], most 3D diffusion models follow a two-stage approach: first learn a latent space of 3D shapes, and then train diffusion in the latent space. However, extending the same success to complex 3D scenes remains a challenge. Such 3D scene modeling faces higher data dimensionality, but with orders of magnitude less training data. Existing object generation methods assume that shapes lie in a canonical space and do not consider the arbitrary-sized nature of 3D scenes, making extending to 3D scenes difficult. In contrast, our approach enables efficient infinite 3D scene generation via training on latent tree scene patches.
3D Scene Generation. Inspired by the success of 2D image generation, various recent methods have proposed reconstructing 3D scenes using 2D generative models [8, 12]. These approaches leverage 2D image generative models to learn to generate consistent novel views; however, designing the camera trajectory and generating a complete 3D scene is challenging due to occlusions. Scene graphs-based 3D generation methods [50, 57] offer an abstract representation for 3D scene generation. Several recent works have introduced diffusion models for 3D scenes [25, 28, 2, 32, 43, 54, 29, 26]. DiffRoom [25], NeuralField-LDM [26], CC3D [2], and XCube [43] focus on bounded generation of 3D scenes, while SSG [28], [32], and SemCity [29] focus on outdoor 3D semantic scenes with limited geometry details. Text-to-3D-based methods [23] typically require hours of optimization. Very recently, BlockFusion [54] proposed to generate larger-scale 3D scenes through latent triplane extrapolation, conditioned on 2D layout inputs. However, 3D scene generation is still in its infancy, and it remains a challenge to efficiently generate unbounded 3D scenes with high-quality, coherent structures, high-frequency details, and plausible layouts. In this work, we propose an efficient unconditional 3D scene generation pipeline that enables seamless infinite 3D scene generation at high quality.
@@ 3.2 Latent Tree Representation for 3D Scenes
Learning. To train the encoders and decoders, our loss measures the $\ell_{2}$ error between the reconstructed TUDFs and the ground truth TUDFs across all the resolution levels:
$\mathcal{L}_{\text{latent}}=\sum\Big{(}L_{i+1}-\mathcal{D}_{i+1}\big{(}\mathcal{E}_{i+1}\left(L_{i+1}\right)\big{)}\Big{)}^{2}.$ (3)
@@ 4 Experiments
Figure 4: Qualitative comparison. We compare unconditional 3D scene generation with diverse 3D diffusion methods PVD [58], NFD [46], and BlockFusion [54]. All methods were trained on houses from the 3D-FRONT dataset [19]. Our latent tree-based 3D scene diffusion approach synthesizes cleaner surfaces with more geometric details and captures diverse furniture objects.
@@
Experimental Setup. We train LT3SD on house-level scene data from the 3D-FRONT dataset [19], a large-scale indoor scene dataset with diverse structures and layouts. Each house contains various room types (e.g., bedrooms, living rooms, kitchens) with different layouts. After filtering out rooms with incorrect furniture scales or no furniture, we retained 6,479 houses. We used an $80\%$/$5\%$/$15\%$ train/validation/test split for both stages of LT3SD and the baselines. Due to open surfaces in scene meshes, consistent signed distance computation is challenging, so we compute UDFs with a voxel size of 0.022m and a truncation value of $\tau=0.1m$. Instead of preprocessing scenes into fixed patches, we continuously sample chunks from 3D scenes during training to reduce overfitting, applying random flips and rotations for data augmentation.
@@
Implementation Details. We implemented our method in Pytorch and ran all experiments on NVIDIA RTX A6000. To first construct latent trees for 3D scenes, we train the encoders and decoders with Adam optimizer [27], using a batch size of $4$ and a learning rate of $1e^{-4}$, which takes from $5$ hours to one day for different resolution levels until convergence. We use $N=3$ levels in our latent tree. For all levels, the latent feature volumes have $4$ channels. We use the Adam optimizer to train the diffusion models, with a batch size of $8$ and a learning rate of $1e^{-4}$, which takes approximately $6$ days until convergence on $2$ GPUs. Note that during the second stage, we infer the latent of each TUDF patch on the fly without storing them on the disk in practice.
@@
Baselines. We compare our method against 3D object diffusion models PVD [58] and NFD [46], as well as the recent 3D scene diffusion approach BlockFusion [54]. PVD is a point-voxel-based diffusion model, while NFD and BlockFusion are triplane-based diffusion models. All baselines were trained on 3D-FRONT. We uniformly sampled 8,192 points on the 3D scene surface for PVD. We used the same TUDF voxel grids for NFD as in our approach. The trained model of BlockFusion is provided by its authors, and we followed the authors’ instructions for comparisons.
@@
Evaluation Metrics. We follow NFD [45] to evaluate generation ability and use an adapted Fréchet Inception Distance (FID) metric based on rendered images of generated meshes to measure perceptual distribution similarity. Following prior work [58, 18, 43], we also assess coverage and diversity using the Coverage (COV), Minimum Matching Distance (MMD), and 1-Nearest Neighbor Accuracy (1-NNA) metrics. COV measures the fraction of test set shapes matched to at least one in the generated set, indicating diversity representation. MMD reflects the average distance from each test shape to its nearest neighbor in the generated set, measuring quality and fidelity. The 1-NNA metric evaluates distribution similarity, with a value closer to $50\%$ indicating statistical indistinguishability between the generated and test distributions.
For FID, we render images of generated and ground-truth 3D meshes by placing cameras at the scene center and rotating them horizontally to capture 20 images per scene. To compute COV, MMD, and 1-NNA, we measure point cloud similarity using Chamfer Distance (CD) and Earth Mover’s Distance (EMD). Points are uniformly sampled from the generated meshes of NFD, BlockFusion, our method, and GT, while the same number of points are randomly sampled from PVD’s generated point clouds.
@@
We show qualitative comparisons in Fig. 4. Our results demonstrate superior surface quality and object detail: PVD generates plausible global structures but produces incomplete geometric information in point clouds; NFD [45] generates plausible scene structures and scales well for complex 3D scenes, but struggles to produce complex arrangements of furniture and smooth surfaces; BlockFusion [54] achieves smooth local surfaces and structural diversity, but struggles in generating meaningful higher-level structures (e.g., Fig. 4-1st row, where walls are misaligned despite improved surface quality). In contrast, our approach generates scenes with both plausible room structures and layouts while maintaining the highest quality in finer-scale details.
@@
Method COV (%) $\uparrow$ MMD ($\times$1e2$\downarrow$ 1-NNA (%) $\downarrow$ FID $\downarrow$
@@
Ours (16 - 128) 28.61 29.98 4.93 30.17 74.51 69.39 59.23
Ours (16 - 32 - 128) 53.10 54.66 3.51 27.30 53.22 52.56 13.39
Tab. 1 shows a quantitative comparison of our LT3SD against baselines. Our approach significantly improves both point-based and perceptual-based metrics, demonstrating its effectiveness in 3D scene generation.
Since the point-cloud-based PVD [58] generates only sparse point clouds, we convert PVD points to triangle meshes with ball-pivoting [4] for FID computation. The triplane-based NFD [45] produces good scene structures, though lacks perceptual quality in local surface smoothness. As shown in Fig. 4, BlockFusion [54] produces recognizable furniture and smooth surfaces, but with quite fragmented room structures (e.g., in Fig. 4, the walls in the 1st and 3rd rows are fragmented and unlikely in structure), resulting in a good FID score but a relatively lower point cloud metrics that rely on more strongly reflecting global structures in the ground-truth distribution.
@@ 4.2 Ablation Study
Latent CM 13.3 M $\times$ 1.00 $\times$ 1.00 4.91e-4
Latent Tree (Ours) 13.3 M $\times$ 0.87 $\times$ 0.80 3.20e-4
Latent Tree Structure. We validate the latent tree structure of our method in Tab. 1 for 3D scene generation. In this ablation, we compare with the one-level latent tree of the same spatial dimension as the coarsest resolution in our latent tree. As shown in Tab. 1-bottom, the multi-level latent tree-based diffusion model outperforms the single-level one. Rather than using a single level of the tree and generating a 3D scene in a single step, our LT3SD adopts a coarse-to-fine strategy for encoding; the diffusion model only generates the higher-frequency component of 3D scenes in each step and, finally, both lower-frequency and higher-frequency components are gracefully combined together and decoded to a high-quality 3D scene. This ablation demonstrates that the hierarchical tree structure is essential for effective generative modeling with diffusion for complex 3D scenes.
Hierarchical Latent Representation. We compare our latent tree with latent Cascaded Models (CMs) (Fig. 2-left), commonly used in 2D image synthesis [22] and 3D generation [32, 43]. Unlike CMs, which independently model redundant information at each level, our latent tree decomposes 3D scenes into coarse geometry and fine latent components, capturing residual high-frequency details. As shown in Tab. 2, with the same autoencoder and parameter count, our model trains more efficiently (requiring less time per level) and achieves higher reconstruction accuracy (lower $l_{2}$ error) on test data. Additionally, its more compact latent representation reduces storage requirements when converting TUDF grids to latent grids. Training time and storage statistics are reported as ratios w.r.t. latent CM; please see the supplemental for more details.
@@ 5 Conclusion
Acknowledgements. This work was supported by the ERC Starting Grant SpatialSem (101076253). Matthias Niessner was supported by the ERC Starting Grant Scan2CAD (804724).
@@ A.1 Intermediate Visualization
We visualize the full inference process of generating a 3D scene unconditionally with LT3SD in Fig. 7. Our method operates patch-wise and coarse-to-fine, as introduced in the main paper. First, starting from the random 3D Gaussian noise on the left of the first row, LT3SD gradually denoises it to the mesh at the end of the row. Then, at the second row, our method autoregressively extrapolates the unknown region in a patch-wise manner with an overlap size of one-half of the patch size until the 3D scene of the specified spatial extent is complete. Finally, in the third row, conditioned at the coarse 3D scene, LT3SD generates the fine details also in a patch-wise manner with the same overlap size. This process is implemented in a batch-wise manner as each patch is only dependent on the same patch of the previous timestep, i.e., all patches at the same timestep are independent of each other. Comparing the coarse-level 3D scene and fine-level 3D scene, we can see that the coarse-level scene determines most of the structure, which ensures a plausible scene layout; the fine-level scene adds rich local detail while keeping the scene structure. The patch-wise and coarse-to-fine 3D scene synthesis enables 3D scene generation with both plausible scene layout and high-quality detail.
A.2 Additional Novelty Analysis Results
@@
A.3 Infinite 3D Scenes
Here, we present more examples of the infinite 3D scene generation beyond Fig. 1 in the main paper. In Fig. 10, we show two more infinite 3D scenes with the resolution of $[4096,2048,128]$ and size of $45.1m\times 90.3m\times 2.8m$. Although trained on the house-level data of 3D-FRONT with only a few connected rooms in each sample, our method generates 3D scenes with diverse room structures, furniture layouts, and varying sizes. We generate infinite 3D scenes in the same manner, i.e., patch-wise and coarse-to-fine. The supplementary video shows more examples of infinite 3D scenes with zoom-ins and intermediate visualization.
A.4 Additional Comparisons
Figure 9: Additional qualitative comparisons. We compare unconditional 3D scene generation with state-of-art 3D diffusion methods PVD [58], NFD [46], and BlockFusion [54]. All methods were trained on houses from the 3D-FRONT dataset [19]. Our latent tree-based 3D scene diffusion approach synthesizes cleaner and more detailed surfaces with diverse furniture objects.
We provide additional comparison results in Fig. 9 using the same model as in the main paper’s qualitative results shown in Fig. 4.
Appendix B Evaluation Metrics

### v2

@@ 2 Related Work
3D Diffusion Models for Shape Generation. Diffusion models have shown unprecedented capability in generating 2D images and videos [47, 6]. Existing 3D diffusion work has focused on synthesizing 3D shapes [7, 20, 49, 60, 16, 33, 15, 22, 11, 55, 58, 30, 14], showing quite promising results. Following LDMs [47], most 3D diffusion models follow a two-stage approach: first learn a latent space of 3D shapes, and then train diffusion in the latent space. However, extending the same success to complex 3D scenes remains a challenge. Such 3D scene modeling faces higher data dimensionality, but with orders of magnitude less training data. Existing object generation methods assume that shapes lie in a canonical space and do not consider the arbitrary-sized nature of 3D scenes, making extending to 3D scenes difficult. In contrast, our approach enables efficient infinite 3D scene generation via training on latent tree scene patches.
3D Scene Generation. Inspired by the success of 2D image generation, various recent methods have proposed reconstructing 3D scenes using 2D generative models [9, 13]. These approaches leverage 2D image generative models to learn to generate consistent novel views; however, designing the camera trajectory and generating a complete 3D scene is challenging due to occlusions. Scene graphs-based 3D generation methods [53, 61] offer an abstract representation for 3D scene generation. Several recent works have introduced diffusion models for 3D scenes [27, 31, 3, 35, 46, 58, 57, 32, 28]. DiffRoom [27], NeuralField-LDM [28], CC3D [3], XCube [46] and TRELLIS [58] focus on bounded generation of 3D scenes, while SSG [31], [35], and SemCity [32] focus on outdoor 3D semantic scene generation, they offer limited geometry detail due to low-resolution data or explicit representation. Text-to-3D-based methods [25] typically require hours of optimization. Very recently, BlockFusion [57] proposed to generate larger-scale 3D scenes through latent triplane extrapolation, conditioned on 2D layout inputs. However, 3D scene generation is still in its infancy, and it remains a challenge to efficiently generate unbounded 3D scenes with high-quality, coherent structures, high-frequency details, and plausible layouts. In this work, we propose an efficient unconditional 3D scene generation pipeline that enables seamless infinite 3D scene generation at high quality.
@@ 3.2 Latent Tree Representation for 3D Scenes
Learning. To train the encoders and decoders, our loss measures the $\ell_{2}$ error between the reconstructed TUDFs and the ground truth TUDFs:
$\mathcal{L}_{\text{latent}}=\Big{(}L_{i+1}-\mathcal{D}_{i+1}\big{(}\mathcal{E}_{i+1}\left(L_{i+1}\right)\big{)}\Big{)}^{2}.$ (3)
@@ 4 Experiments
Figure 4: Qualitative comparison. We compare unconditional 3D scene generation with diverse 3D diffusion methods PVD [62], NFD [49], BlockFusion [57], and XCube [46]. All methods were trained on houses from the 3D-FRONT dataset [21]. Our latent tree-based 3D scene diffusion approach synthesizes cleaner surfaces with more geometric details and captures diverse furniture objects.
@@
Experimental Setup. We train LT3SD on house-level scene data from the 3D-FRONT dataset [21], a large-scale indoor scene dataset with diverse structures and layouts. Each house contains various room types (e.g., bedrooms, living rooms, kitchens) with different layouts. After filtering out rooms with incorrect furniture scales or no furniture, we retained 6,479 houses. We used an $80\%$/$5\%$/$15\%$ train/validation/test split for both stages of LT3SD and the baselines. Due to open surfaces in scene meshes, consistent signed distance computation is challenging, so we compute UDFs with a voxel size of $2.2cm$ and a truncation value of $\tau=10cm$. Instead of preprocessing scenes into fixed patches, we continuously sample chunks from 3D scenes during training to reduce overfitting, applying random flips and rotations for data augmentation.
@@
Implementation Details. We implemented our method in Pytorch and ran all experiments on NVIDIA RTX A6000. To first construct latent trees for 3D scenes, we train the encoders and decoders with Adam optimizer [29], using a batch size of $4$ and a learning rate of $10^{-4}$, which takes from $5$ hours to one day for different resolution levels until convergence. We use $N=3$ levels in our latent tree. For all levels, the latent feature volumes have $4$ channels. We use the Adam optimizer to train the diffusion models, with a batch size of $8$ and a learning rate of $10^{-4}$, which takes approximately $6$ days until convergence on $2$ GPUs. Note that during the second stage, we infer the latent of each TUDF patch on the fly without storing them on the disk in practice.
@@
Baselines. We compare our method against 3D object diffusion models PVD [62] and NFD [49], as well as the recent 3D scene diffusion approach BlockFusion [57] and XCube [46]. PVD is based on pointclouds, XCube is based on latent sparse voxels, while NFD and BlockFusion are based on triplanes. All baselines were trained on 3D-FRONT. We uniformly sampled 8,192 points on the 3D scene surface for PVD. We used the same TUDF voxel grids for NFD and XCube as in our approach. The trained model of BlockFusion is provided by its authors, and we followed the authors’ instructions for comparisons.
@@
Evaluation Metrics. We follow NFD [48] to evaluate generation ability and use an adapted Fréchet Inception Distance (FID) metric based on rendered images of generated meshes to measure perceptual distribution similarity. We render images of generated and ground-truth 3D meshes by placing cameras at the scene center and rotating them horizontally to capture 20 images per scene. Following prior work [62, 20, 46], we also assess coverage and diversity using the Coverage (COV), Minimum Matching Distance (MMD), and 1-Nearest Neighbor Accuracy (1-NNA) metrics. COV measures the fraction of test set shapes matched to at least one in the generated set, indicating diversity representation. MMD reflects the average distance from each test shape to its nearest neighbor in the generated set, measuring quality and fidelity. The 1-NNA metric evaluates distribution similarity, with a value closer to $50\%$ indicating statistical indistinguishability between the generated and test distributions. We measure point cloud similarity using Chamfer Distance (CD) and Earth Mover’s Distance (EMD). Points are uniformly sampled from the generated meshes of NFD, BlockFusion, our method, and GT, while the same number of points are randomly sampled from PVD’s generated point clouds. Please refer to the supplemental for more details.
@@
We show qualitative comparisons in Fig. 4. Our results demonstrate superior surface quality and object detail: PVD generates plausible global structures but produces incomplete geometric information in point clouds; NFD [48] generates plausible scene structures and scales well for complex 3D scenes, but struggles to produce complex arrangements of furniture and smooth surfaces; BlockFusion [57] achieves smooth local surfaces and structural diversity, but struggles in generating meaningful higher-level structures (e.g., Fig. 4-1st row, where walls are misaligned despite improved surface quality). XCube [46] produces high-quality scene structures but struggles with local details, like wrinkles on the bed In contrast, our approach generates scenes with both plausible room structures and layouts while maintaining the highest quality in finer-scale details.
@@
Method COV (%) $\uparrow$ MMD (×$\times 10^{-2}$$\downarrow$ 1-NNA (%) $\downarrow$ FID $\downarrow$
@@
XCube [62] 48.60 47.71 3.35 26.14 56.45 59.59 55.35
Ours (17.6 - 2.2) cm 28.61 29.98 4.93 30.17 74.51 69.39 59.23
Ours (8.8 - 2.2) cm 40.87 44.21 4.90 31.50 84.34 81.57 50.07
Ours (17.6 - 8.8 - 2.2) cm 53.10 54.66 3.51 27.30 53.22 52.56 13.39
Tab. 1 shows a quantitative comparison of our LT3SD against baselines. Our approach significantly improves both point-based and perceptual-based metrics, demonstrating its effectiveness in 3D scene generation. Since the point-cloud-based PVD [62] generates only sparse point clouds, we convert PVD points to triangle meshes with ball-pivoting [5] for FID computation. The triplane-based NFD [48] produces good scene structures, though lacks perceptual quality in local surface smoothness. As shown in Fig. 4, BlockFusion [57] produces recognizable furniture and smooth surfaces, but with quite fragmented room structures (e.g., in Fig. 4, the walls in the 1st and 3rd rows are fragmented and unlikely in structure), resulting in a good FID score but a relatively lower point cloud metrics that rely on more strongly reflecting global structures in the ground-truth distribution. XCube [46] produces compelling 3D scene structure but lacks accurate details by generating complex 3D scenes in a single step without decomposing them to patches.
@@ 4.2 Ablation Study
Latent CM 13.3 M $\times$ 1.00 $\times$ 1.00 4.91 $\times$ $10^{-4}$
Latent Tree (Ours) 13.3 M $\times$ 0.87 $\times$ 0.80 3.20 $\times$ $10^{-4}$
Latent Tree Structure. We validate the latent tree structure of our method in Tab. 1 for 3D scene generation. In this ablation, we compare our full model with latent trees containing only one level. As shown in Tab. 1-bottom, the multi-level latent tree-based diffusion model outperforms the single-level ones. Rather than using a single level of the tree and generating a 3D scene in a single step, our LT3SD adopts a coarse-to-fine strategy for encoding; the diffusion model only generates the higher-frequency component of 3D scenes in each step and, finally, both lower-frequency and higher-frequency components are gracefully combined together and decoded to a high-quality 3D scene. This ablation demonstrates that the hierarchical tree structure is essential for effective generative modeling with diffusion for complex 3D scenes.
Hierarchical Latent Representation. We compare our latent tree with latent Cascaded Models (CMs) (Fig. 2-left), commonly used in 2D image synthesis [24] and 3D generation [35, 46]. Unlike CMs, which independently model redundant information at each level, our latent tree decomposes 3D scenes into coarse geometry and fine latent components, capturing residual high-frequency details. As shown in Tab. 2, with the same autoencoder and parameter count, our model trains more efficiently (requiring less time per level) and achieves higher reconstruction accuracy (lower $l_{2}$ error) on test data. Additionally, its more compact latent representation (fewer latent channels) reduces storage requirements when converting TUDF grids to latent grids. Training time and storage statistics are reported as ratios w.r.t. latent CM; please see the supplemental for more details.
@@ 5 Conclusion
Acknowledgements. This work was supported by the ERC Starting Grant SpatialSem (101076253). Matthias Nießner was supported by the ERC Starting Grant Scan2CAD (804724). We would like to thank Xuanchi Ren for helpful suggestions on training XCube.
@@ A.1 Intermediate Visualization
We visualize the full inference process of generating a 3D scene unconditionally with LT3SD in Fig. 7. Our method operates patch-wise and coarse-to-fine, as introduced in the main paper. First, starting from the random 3D Gaussian noise on the left of the first row, LT3SD gradually denoises it to the mesh at the end of the row. Then, at the second row, our method autoregressively extrapolates the unknown region in a patch-wise manner with an overlap size of one-half of the patch size, which empirically provides sufficient context and reduces seams until the 3D scene of the specified spatial extent is complete. Finally, in the third row, conditioned at the coarse 3D scene, LT3SD generates the fine details also patch-wise with the same overlap size. This process is implemented batch-wise as each patch is only dependent on the same patch of the previous timestep, i.e., all patches at the same timestep are independent of each other. Comparing the coarse-level 3D scene and fine-level 3D scene, we can see that the coarse-level scene determines most of the structure, which ensures a plausible scene layout; the fine-level scene adds rich local detail while keeping the scene structure. The patch-wise and coarse-to-fine 3D scene synthesis enables 3D scene generation with both plausible scene layout and high-quality detail.
A.2 Additional Infinite 3D Scenes
Unlike retrieval-based methods (e.g., DiffuScene [53]), focusing on object layouts for single rooms (without walls/structures), our approach synthesizes large-scale scenes, including objects/walls/floors, while maintaining coherent structures at the room scale. LT3SD extends infinitely through outpainting, making it well-suited for open-world games, large-scale robotic training, and film asset creation–a step toward scalable indoor scene generation.
Here, we present more examples of the infinite 3D scene generation beyond Fig. 1 in the main paper. In Fig. 11, we show two more infinite 3D scenes with the resolution of $[4096,2048,128]$ and size of $90.3m\times 45.1m\times 2.8m$. Although trained on the house-level data of 3D-FRONT with only a few connected rooms in each sample, our method generates 3D scenes with diverse room structures, furniture layouts, and varying sizes. We generate infinite 3D scenes in the same manner, i.e., patch-wise and coarse-to-fine. The supplementary video shows more examples of infinite 3D scenes with zoom-ins and intermediate visualization.
A.3 Additional Novelty Analysis Results
@@
A.4 Outdoor 3D Scene Generation
Figure 9: Outdoor 3D Scene Generation.
In this section, we provide 3D outdoor scene generation visualization in Fig. 9. Unlike methods that assume a specific 3D scene structure, our approach employs a unified voxel-based representation for 3D scenes, allowing seamless extension beyond indoor scenes to diverse environments such as outdoor 3D scenes. As shown in Fig. 9, we train our models on the 3D city asset [1] in a patch-wise manner, as described in the main paper, and generate large 3D scenes with a resolution of $[1536,1536,512]$, where $512$ is the vertical dimension.
Appendix B Evaluation
Figure 10: Additional qualitative comparisons. We compare unconditional 3D scene generation with state-of-art 3D diffusion methods PVD [62], NFD [49], BlockFusion [57], and XCube [46]. All methods were trained on houses from the 3D-FRONT dataset [21]. Our latent tree-based 3D scene diffusion approach synthesizes cleaner and more detailed surfaces with diverse furniture objects.
B.1 Metrics
For FID, we set the patch size to be $5.6m\times 5.6m\times 2.8m$ to consider both geometry quality and scene layout. For point clouds-based metrics, we split each generated and ground-truth patch into four smaller ones with the size of $2.8m\times 2.8m\times 2.8m$ to make sure complex structures are represented with a limited number of sampled points.